OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 15 miles per gallon of gas and the second has a fuel efficiency of 30 miles per gallon of gas. During one particular week, the two cars went a combined total of 1725 miles, for a total gas consumption of 75 gallons. How many gallons were consumed by each of the two cars that week? First car: Second car: I truly loathe word problems...please help? Thank you!

OpenStudy (anonymous):

Hikx2bay thanks for looking at my problem :)

OpenStudy (anonymous):

no worries buddy 1st car: 575miles 2nd car:1150 miles please dont ask me how i got these answers ;)...guess work...

OpenStudy (anonymous):

As you wish...thank you

OpenStudy (anonymous):

actually there's some logic behind it, let me try explain

OpenStudy (anonymous):

car 1 does x miles / gallon car 2 does 2.x miles / gallon

OpenStudy (anonymous):

so you have x + 2x = 3x miles / gallon and use that proportion on the total distance 1725 miles

OpenStudy (anonymous):

1725 * 1/3 = 575 miles for car 1 1725 * 2/3 = 1150 miles for car 2

OpenStudy (anonymous):

how does that sound?

OpenStudy (anonymous):

like it makes sense...whoah

OpenStudy (anonymous):

:)